# Learned feature representations are biased by complexity, learning order, position, and more

### v1

@@ B.11 Low-dimensional visualizations
Figure 39: Principal components visualizations of penultimate representations from MLPs that are trained to output a fixed hard feature (sum % 2), together with different numbers of easy linear features (increasing from 1-4 across rows). As easy features are added, their labels (colors) spread across more principal components (columns), and thus the dimension along which the hard features (shapes) separate is pushed into higher and higher principal components. With 1 easy feature (top row), the first two PCs are clustered according to the easy feature (39(a)), but the hard feature is clustered along the third principal component (39(b)). When two easy features are learned, the third principal component is used to help represent them, and the hard feature is pushed into the fourth principal component (39(e)). When three or four easy features are learned, the hard feature is pushed into the 5th or 6th PCs, respectively (39(i) & 39(l)). There is a remarkably consistent progression that as each easy feature is added, another PC is captured, and the hard feature is pushed one PC higher.

### v2

@@ B.12 Low-dimensional visualizations
Figure 41: Principal components visualizations of penultimate representations from MLPs that are trained to output a fixed hard feature (sum % 2), together with different numbers of easy linear features (increasing from 1-4 across rows). As easy features are added, their labels (colors) spread across more principal components (columns), and thus the dimension along which the hard features (shapes) separate is pushed into higher and higher principal components. With 1 easy feature (top row), the first two PCs are clustered according to the easy feature (41(a)), but the hard feature is clustered along the third principal component (41(b)). When two easy features are learned, the third principal component is used to help represent them, and the hard feature is pushed into the fourth principal component (41(e)). When three or four easy features are learned, the hard feature is pushed into the 5th or 6th PCs, respectively (41(i) & 41(l)). There is a remarkably consistent progression that as each easy feature is added, another PC is captured, and the hard feature is pushed one PC higher.